# Rigorously Characterizing Dynamics with Machine Learning

### v1

@@ 4.1. Simplicial Complexes
A collection $\mathcal{U}$ of $d$-dimensional simplicies (with boundary) is a good closed cover of $X$ if ${\mathcal{U}}^{0}:=\left\{{\mathop{\mathrm{int}}\nolimits(U)}\,\left|\,{U\in{\mathcal{U}}}\right.\right\}$ is an open cover of $X$ and the intersection of any finite subcollection of ${\mathcal{U}}$ is either empty or acyclic. Let $\text{sd}^{n}({\mathcal{X}})$ denote the simplicial complex obtained by performing $n$ barycentric subdivisions to ${\mathcal{X}}$.
@@
Let ${\mathcal{X}}$ denote a finite simplicial covering of $X$. There exists $k\in{\mathbb{N}}$ such that the collection ${\mathcal{U}}^{o}$ of all open stars of vertices of $\text{sd}^{k}({\mathcal{X}})$ is a finite open cover of $X$. Furthermore, ${\mathcal{U}}:=\left\{{\mathop{\mathrm{cl}}\nolimits(U)}\,\left|\,{U\in{\mathcal{U}}^{o}}\right.\right\}$ is a finite good closed cover of $X$.
@@
That ${\mathcal{U}}$ is a good open cover of $X$ is straightforward. Whether the closures form a good closed cover depends on the choice of triangulation ${\mathcal{T}}$. A straightforward, but technical argument, shows that if one starts with ${\mathcal{T}}^{(2)}$, the second barycentric subdivision of ${\mathcal{T}}$, then ${\mathcal{U}}$ has the desired property [35]. ∎
@@ 4.2. Outer Approximations
We use a combinatorial multivalued map ${\mathcal{F}}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ defined on a geometric simplicial complex ${\mathcal{X}}$ to encode dynamics and set
@@
${\mathcal{F}}_{\rho}(\xi):=\left\{{\eta\in{\mathcal{X}}}\,\left|\,{B_{\rho}(f(|\xi|))\cap|\eta|\neq\emptyset}\right.\right\},$
@@
[17, Corollary 2.6] A combinatorial multivalued map ${\mathcal{F}}$ is an outer approximation of $f$ if and only if ${\mathcal{F}}$ encloses the minimal combinatorial outer approximation ${\mathcal{F}}_{0}$ of $f$.
@@
The following result is a special case of [17, Proposition 7.3].
@@
Applying Proposition 2.17 we obtain the following corollary.
@@
Let ${\mathcal{U}}$ be a good closed cover of $X$, let ${\mathcal{X}}$ be a simplicial cover of $X$, and assume that ${\mathcal{F}}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ is a combinatorial multivalued map inscribed into ${\mathcal{U}}$. Assume that ${\mathcal{N}}_{0}\subset{\mathcal{N}}_{1}$ are subcomplexes of ${\mathcal{X}}$ with the property that ${\mathcal{F}}({\mathcal{N}}_{i})\subset{\mathcal{N}}_{i}$. Then, as is shown in [13] minor modifications of the algorithms presented in [14] give rise to a well defined homomorphism
@@
We leave it to the reader to check that ${\mathcal{N}}(q)\in{\mathsf{J}}^{\vee}({\mathsf{Invset}}^{+}({\mathcal{F}}))$ (see [19]). Thus, $(|{\mathcal{N}}(q)|,|{\mathcal{N}}(q)^{<}|)$ is an index pair and the Conley index of $\mathop{\mathrm{Inv}}\nolimits(|{\mathcal{N}}(q)\setminus{\mathcal{N}}(q)^{<}|,f)$ is given by the shift equivalence class of ${\mathcal{F}}_{*}\colon H_{*}({\mathcal{N}}(q),{\mathcal{N}}(q)^{<})\to H_{*}({\mathcal{N}}(q),{\mathcal{N}}(q)^{<})$.
@@
4.3. Piecewise linear approximations
@@
Throughout this section, $G$ denotes a function of the form described in Theorem 1.2. Given $G$ and ${\mathcal{P}}$, let ${\mathcal{P}}_{X}:=\left\{{\zeta\cap X}\,\left|\,{\zeta\in{\mathcal{P}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \zeta\cap X\neq\emptyset}\right.\right\}$. For simplicity of exposition we assume each polytope $\zeta\in{\mathcal{P}}_{X}$ is subdivided into a finite set of simplicies, giving rise to a finite simplicial complex ${\mathcal{X}}$. For the remainder of this paper we use ${\mathcal{G}}_{\rho}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ to denote the $\rho$-minimal combinatorial outer approximation of $G$ on ${\mathcal{X}}$.
@@
Let $f\colon X\to X$ be a continuous function on a compact subset of ${\mathbb{R}}^{d}$. Let $N\subset{\mathsf{ABlock}}(f)$. The tolerance of $N$ is defined to be
$\tau(N):=\sup\left\{{\nu>0}\,\left|\,{B_{\nu}(f(N))\subset N}\right.\right\}$
The assumption that $f$ is continuous implies that given $N\subset{\mathsf{ABlock}}(f)$ there exists $a(N)>0$ such that
@@
Consider $G\colon X\to X$ such that $\sup_{x\in X}\|f(x)-G(x)\|<\rho(N)$ and assume $\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho(N)$. Let $s(N):=\left\{{\xi\in{\mathcal{X}}}\,\left|\,{|\xi|\cap N\neq\emptyset}\right.\right\}$ and let $\mathop{\mathrm{bdy}}\nolimits_{\mathcal{X}}(N):=\left\{{\xi\in{\mathcal{X}}}\,\left|\,{|\xi|\cap\mathop{\mathrm{bdy}}\nolimits(N)\neq\emptyset}\right.\right\}$. Then
@@
(i) Recall that ${\mathcal{G}}_{\rho(N)}$ approximates $G$. By definition of $\rho(N)$,
${\mathcal{G}}_{\rho(N)}(s(N))\subset B_{2\rho(N)}(f(N))\subsetneq B_{\tau(N)/2}(f(N)).$
@@
The assumption $\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho(N)$ can always be satisfied by taking barycentric subdivisions of the original simplicial complex over which $G$ is defined [15, pp. 120].
@@
Let $G\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ be a fully-connected feedforward neural network with piecewise linear activation on the simplicial complex ${\mathcal{X}}$. Fix $\epsilon>0$. Then, there exists $n=n(\epsilon)$ such that
@@
Let $L$ denote the Lipschitz constant of $G$. Then $\max_{\xi\in\text{sd}^{n}({\mathcal{X}})}\mathop{\mathrm{diam}}\nolimits(G(|\xi|))\leq L\mathop{\mathrm{diam}}\nolimits(\text{sd}^{n}({\mathcal{X}}))$. By [15, pp. 120],
@@
Figure 2. (a) $X$ is denoted by the thick black line. Morse representation ${\mathsf{MR}}=\left\{{M_{0},M_{1},M_{2}}\right\}$ with partial order $M_{0}<_{{\mathsf{MR}}_{f}}M_{2}$ and $M_{1}<_{{\mathsf{MR}}_{f}}M_{2}$. (b) $N_{0}=k(\psi(M_{0}))$, $N_{1}=k(\psi(M_{1}))$ and $N_{2}=k(\psi(M_{2}))=X$ (not shown) are join irreducible attracting blocks. (c) Hash marks represent the grid ${\mathcal{X}}$ with diameter less than $\rho({\mathsf{N}})$ that is used to define $s(N_{0})$ and $s(N_{1})$. (d) The Morse graph for ${\mathcal{G}}$ is ${\mathsf{MG}}({\mathcal{G}})=\left\{{q_{i}}\,\left|\,{i=0,\ldots,7}\right.\right\}$ with partial order $q_{0}<_{{\mathsf{MG}}({\mathcal{G}})}q_{4}<_{{\mathsf{MG}}({\mathcal{G}})}q_{5}<_{{\mathsf{MG}}({\mathcal{G}})}q_{7}$, $q_{1}<_{{\mathsf{MG}}({\mathcal{G}})}q_{6}<_{{\mathsf{MG}}({\mathcal{G}})}q_{7}$, $q_{1}<_{{\mathsf{MG}}({\mathcal{G}})}q_{3}$, and $q_{2}<_{{\mathsf{MG}}({\mathcal{G}})}q_{3}$.
By Remark 2.5 there exists a unique finite lattice of attractors ${\mathsf{A}}$ for which there is a poset isomorphism $\psi\colon{\mathsf{MR}}\to{\mathsf{J}}^{\vee}({\mathsf{A}})$. By [19, pp. 1179], [19, Theorem 1.2], and Proposition 4.6 there exists a sublattice ${\mathsf{N}}$ of ${\mathsf{ABlock}}(f)$ (with minimal element ${\mathbf{0}}_{\mathsf{N}}:=\emptyset$ and maximal element ${\mathbf{1}}_{\mathsf{N}}=X$) and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$ for all $A\in{\mathsf{A}}$. Figure 2(B) shows a choice of attracting blocks $N_{i}=k(\psi(M_{i}))$ for $i=0,1$. By [19, Theorem 1.2], $k(\psi(M_{2}))=X$.
While [19, Theorem 1.2] provides for the existence of a sublattice ${\mathsf{N}}$, there is no claim that it is unique (see Example 2.15). Therefore, we emphasize that for the remainder of this section we assume that ${\mathsf{N}}$ is fixed with minimal element ${\mathbf{0}}_{\mathsf{N}}:=\emptyset$ and maximal element ${\mathbf{1}}_{\mathsf{N}}=X$. For each $N\in{\mathsf{N}}$, choose $\rho(N)>0$ satisfying Proposition 4.12. Set
$\rho({\mathsf{N}})=\min\left\{{\rho(N)}\,\left|\,{N\in{\mathsf{N}}}\right.\right\}.$
For the remainder of this section, we consider $G\colon X\to X$ such that
$\sup_{x\in X}\|f(x)-G(x)\|<\rho({\mathsf{N}})\quad\text{and}\quad\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho({\mathsf{N}}).$
Given $N\in{\mathsf{N}}$, define $s(N):=\left\{{\xi\in{\mathcal{X}}}\,\left|\,{\mathop{\mathrm{cl}}\nolimits(\xi)\cap N\neq\emptyset}\right.\right\}$ and set ${\mathsf{K}}:=\left\{{s(N)}\,\left|\,{N\in{\mathsf{N}}}\right.\right\}$. The intervals defined by the hash marks in Figure 2(C) are the polytopes for the complex ${\mathcal{X}}$ with $\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho({\mathsf{N}})$. The sets $s(N_{i})$, for $i=0,1$ are also shown.
@@
We now show that $M\cap k(\psi(M)^{<})=\emptyset$. By [18, Proposition 2.11],
$\omega(M\cap k(\psi(M)^{<}),f)\subset\omega(M,f)\cap\omega(k(\psi(M)^{<}),f).$
Then
@@
By [18, Proposition 2.11], $M\cap k(\psi(M)^{<})=\emptyset$.
@@
Let $\pi\colon{\mathcal{X}}\to\bar{{\mathcal{X}}}$ denote the quotient map that reduces ${\mathcal{G}}$ to its condensation graph $\bar{{\mathcal{G}}}$. The intervals highlighted in Figure 2(D) are the Morse nodes of the Morse graph ${\mathsf{MG}}({\mathcal{G}})=\left\{{q_{j}}\,\left|\,{j=0,\ldots,7}\right.\right\}$ of the outer approximation ${\mathcal{G}}$ of $G$. We use this hypothetical example to highlight two challenges that arise from using the dynamics computed via ${\mathcal{G}}$ to describe the dynamics of $f$: spurious recurrence and overly refined recurrence.
Before turning to the challenges we note that the Morse nodes $q_{0}$ and $q_{7}$ identify the location of the Morse sets $M_{0}$ and $M_{2}$ of $f$, respectively, in that $M_{0}\subset\pi^{-1}(q_{0})$ and $M_{0}\subset\pi^{-1}(q_{2})$. Furthermore, as is shown in the proof of Theorem 1.4, $\mathop{\mathrm{Con}_{*}}\nolimits(M_{0})\sim\mathop{\mathrm{Con}_{*}}\nolimits(q_{0})$ and $\mathop{\mathrm{Con}_{*}}\nolimits(M_{2})\sim\mathop{\mathrm{Con}_{*}}\nolimits(q_{7})$. This represents the ideal information that can be provided by ${\mathcal{G}}$.
@@
Overly refined recurrence arises when $G$ provides more information about the dynamics of $f$ than is asked for in Theorem 1.4. Stated differently, there exists a Morse representation of $f$ that is finer than ${\mathsf{MR}}$, and ${\mathcal{G}}$ can identify this finer Morse representation. Therefore, to satisfy the conclusion of Theorem 1.4 we collapse this information to construct the desired Morse representation.
@@
$I(M):=\left\{{q\in{\mathsf{MG}}({\mathcal{G}})}\,\left|\,{\pi^{-1}(q)\subset s(k(\psi(M)))\setminus s(k(\psi(M)^{<}))}\right.\right\}.$
@@
By Proposition 5.3, there exists $q\in{\mathsf{MG}}({\mathcal{G}})$ such that $M\cap|\pi^{-1}(q)|\neq\emptyset$. By Proposition 5.2, $M\subset|s(k(\psi(M)))\setminus s(k(\psi(M)^{<}))|$. This implies that $\pi^{-1}(q)\cap s(k(\psi(M)))\neq\emptyset$. By Lemma 5.4, $\pi^{-1}(q)\subset s(k(\psi(M)))$. Then $\pi^{-1}(q)\cap s(k(\psi(M)^{<}))=\emptyset$, for otherwise Lemma 5.4 implies that $\pi^{-1}(q)\subset s(k(\psi(M)^{<}))$, contradicting the fact that $M\cap|\pi^{-1}(q)|\neq\emptyset$ and $M\subset|s(k(\psi(M)))\setminus s(k(\psi(M)^{<}))|$. Therefore, $I(M)\neq\emptyset$.
If $I(M)$ consists of a unique element, then the result is trivially true.
Thus, we need to show that if $q_{0},q_{1}\in I(M)$ and $q_{0}<_{{\mathsf{MG}}({\mathcal{G}})}q^{\prime}<_{{\mathsf{MG}}({\mathcal{G}})}q_{1}$, then $q^{\prime}\in I(M)$. We provide a proof by contradiction, i.e., we assume that $q^{\prime}\not\in I(M)$.
@@
$\left\{{I(M)}\,\left|\,{M\in{\mathsf{MR}}}\right.\right\}$ is a partition of ${\mathsf{MG}}({\mathcal{G}})$.
@@
Recall that $\psi:{\mathsf{MR}}\to{\mathsf{J}}^{\vee}({\mathsf{A}})$ is a poset isomorphism and $k({\mathsf{J}}^{\vee}({\mathsf{A}}))={\mathsf{J}}^{\vee}({\mathsf{N}})$. By [6, Proposition 2.45], $X={\mathbf{1}}_{\mathsf{N}}=\bigvee_{M\in{\mathsf{MR}}}k(\psi(M))=\bigcup_{M\in{\mathsf{MR}}}k(\psi(M))$. Therefore, ${\mathcal{X}}=\bigcup_{M\in{\mathsf{MR}}}s(k(\psi(M)))$.
@@
So assume $\psi(M_{0})^{<}\not\in{\mathsf{J}}({\mathsf{A}})$. As given by Proposition 5.1 and [19, Theorem 1.1], $s$ and $k$ are isomorphisms, which implies $\psi(M_{0})^{<}\neq\emptyset$. Then by [6, Proposition 2.45], $\psi(M_{0})^{<}$ can be written as the non-empty join $\psi(M_{0})^{<}=\bigvee_{M<M_{0}}s(k(\psi(M)))$. Thus, $\pi^{-1}(q)\cap s(k(\psi(M_{1})))\neq\emptyset$ for some $M_{1}<M_{0}$ and we repeat the argument.
These arguments can only be repeated finitely many times and hence there exists $M\in{\mathsf{MR}}$ such that $q\in I(M)$.
By definition, $q$ belongs to a unique $I(M)$, and therefore, $\left\{{I(M)}\,\left|\,{M\in{\mathsf{MR}}}\right.\right\}$ is a partition of ${\mathsf{MG}}({\mathcal{G}})$. ∎

### v2

@@ 4.1. Simplicial Complexes
[…]
@@
[…]
@@
That $\mathcal{U}$} is a good open cover of $X$ is straightforward. Whether the closures form a good closed cover depends on the choice of triangulation $\mathcal{T}$}. A straightforward, but technical argument, shows that if one starts with ${\mathcal{T}}^{(2)}$, the second barycentric subdivision of $\mathcal{T}$}, then $\mathcal{U}$} has the desired property [ramras]. ∎
@@ 4.2. Outer Approximations
We use a combinatorial multivalued map
${\mathcal{F}}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ defined on a geometric simplicial complex $\mathcal{X}$} to encode dynamics and set
@@
[…]
@@
[kalies:mischaikow:vandervorst:05, Corollary 2.6] A combinatorial multivalued map $\mathcal{F}$} is an outer approximation of $f$ if and only if $\mathcal{F}$} encloses the minimal combinatorial outer approximation ${\mathcal{F}}_{0}$ of $f$.
@@
The following result is a special case of [kalies:mischaikow:vandervorst:05, Proposition 7.3].
@@
Proposition 4.6 together with [kalies:mischaikow:vandervorst:15, Theorem 1.2] [^1] The remarks of [kalies:mischaikow:vandervorst:15, pp. 1179] guarantee that the assumptions of the theorem can be satisfied. implies the following result.
Proposition 4.7.
Let ${\mathsf{A}}\subset{\mathsf{Att}}(f)$ be a finite lattice of attractors for $f$. There exists a sublattice $\mathsf{N}$} of ${\mathsf{ABlock}}(f)$ and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$ for all $A\in{\mathsf{A}}$.
From Proposition 2.17 and Proposition 4.6 we obtain the following corollary.
@@
Let $\mathcal{U}$} be a good closed cover of $X$, let $\mathcal{X}$} be a simplicial cover of $X$, and assume that ${\mathcal{F}}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ is a combinatorial multivalued map inscribed into $\mathcal{U}$}. Assume that ${\mathcal{N}}_{0}\subset{\mathcal{N}}_{1}$ are subcomplexes of $\mathcal{X}$} with the property that ${\mathcal{F}}({\mathcal{N}}_{i})\subset{\mathcal{N}}_{i}$. Then, as is shown in [harker:kokubu:mischaikow:pilarczyk] minor modifications of the algorithms presented in [harker:mischaikow:mrozek:nanda] give rise to a well defined homomorphism
@@
We leave it to the reader to check that ${\mathcal{N}}(q)\in{\mathsf{J}}^{\vee}({\mathsf{Invset}}^{+}({\mathcal{F}}))$ (see [kalies:mischaikow:vandervorst:15]). Thus, $(|{\mathcal{N}}(q)|,|{\mathcal{N}}(q)^{<}|)$ is an index pair and the Conley index of $\mathop{\mathrm{Inv}}\nolimits(|{\mathcal{N}}(q)\setminus{\mathcal{N}}(q)^{<}|,f)$ is given by the shift equivalence class of ${\mathcal{F}}_{*}\colon H_{*}({\mathcal{N}}(q),{\mathcal{N}}(q)^{<})\to H_{*}({\mathcal{N}}(q),{\mathcal{N}}(q)^{<})$.
@@
4.3. Approximating attracting blocks
@@
Throughout this section, $G$ denotes a function satisfying P1 - P3 and $\mathcal{X}$} is a simplicial complex satisfying X1 and X2. For the remainder of this paper we use ${\mathcal{G}}_{\rho}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ to denote the $\rho$-minimal combinatorial outer approximation of $G$ on $\mathcal{X}$}. Note that P2 and P3 allow for (at least conceptually) the computation of ${\mathcal{G}}_{\rho}$.
We remark that if $G$ satisfies Theorem 1.5 then $\mathcal{X}$} can be derived from $\mathcal{P}$} by subdividing each polytope into a finite set of simplicies which in turn simplifies the evaluation of ${\mathcal{G}}_{\rho}$.
@@
Let $f\colon X\to X$ be a continuous function on a compact subset of ${\mathbb{R}}^{d}$. Let $N\in{\mathsf{ABlock}}(f)$. The tolerance of $N$ is defined to be
[…]
The assumption that $f$ is continuous implies that given $N\in{\mathsf{ABlock}}(f)$ there exists $a(N)>0$ such that
@@
[…]
@@
(i) Recall that ${\mathcal{G}}_{\rho(N)}$ approximates $G$. Since $\sup_{x\in X}\|f(x)-G(x)\|<\rho(N)$ and $\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho(N)$,
${\mathcal{G}}_{\rho(N)}(s(N))\subset B_{2\rho(N)}(G(B_{\rho(N)}(N)))\subset B_{3\rho(N)}(f(B_{\rho(N)}(N))).$
By definition of $\rho(N)$,
$B_{3\rho(N)}(f(B_{\rho(N)}(N)))\subset B_{3\rho(N)}(B_{\tau(N)/4}(f(N)))\subsetneq B_{\tau(N)}(f(N)).$
@@
The assumption $\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho(N)$ can always be satisfied by taking barycentric subdivisions of the original simplicial complex over which $G$ is defined [Hatcher, pp. 120].
@@
Let $\epsilon>0$. Then, there exists $n=n(\epsilon)$ such that
@@
Let $L$ denote the Lipschitz constant of $G$. Then
$\max_{\xi\in\text{sd}^{n}({\mathcal{X}})}\mathop{\mathrm{diam}}\nolimits(G(|\xi|))\leq L\mathop{\mathrm{diam}}\nolimits(\text{sd}^{n}({\mathcal{X}}))$. By [Hatcher, pp. 120],
@@
[…]
By Remark 2.5 there exists a unique finite lattice of attractors $\mathsf{A}$} for which there is a poset isomorphism $\psi\colon{\mathsf{MR}}\to{\mathsf{J}}^{\vee}({\mathsf{A}})$.
By Proposition 4.7 there exists a sublattice $\mathsf{N}$} of ${\mathsf{ABlock}}(f)$ (with minimal element ${\mathbf{0}}_{\mathsf{N}}:=\emptyset$ and maximal element ${\mathbf{1}}_{\mathsf{N}}=X$) and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$ for all $A\in{\mathsf{A}}$. Figure 3(B) shows a choice of attracting blocks $N_{i}=k(\psi(M_{i}))$ for $i=0,1$. By [kalies:mischaikow:vandervorst:15, Theorem 1.2], $k(\psi(M_{2}))=X$.
While [kalies:mischaikow:vandervorst:15, Theorem 1.2] provides for the existence of a sublattice $\mathsf{N}$}, there is no claim that it is unique (see Example 2.15). Therefore, we emphasize that for the remainder of this section we assume that $\mathsf{N}$} is fixed with minimal element ${\mathbf{0}}_{\mathsf{N}}:=\emptyset$ and maximal element ${\mathbf{1}}_{\mathsf{N}}=X$. For each $N\in{\mathsf{N}}$, choose $\rho(N)>0$ satisfying Proposition 4.13. Set
[…]
For the remainder of this section, we consider $G\colon X\to X$ satisfying P1 where $\mathop{\mathrm{diam}}\nolimits({\mathcal{X}})<\rho({\mathsf{N}})$ and set ${\mathcal{G}}={\mathcal{G}}_{\rho({\mathsf{N}})}$.
[…]
@@
We now show that $M\cap k(\psi(M)^{<})=\emptyset$. We leave it to the reader to check that given $U,V\subset X$, $\omega(U\cap V,f)\subset\omega(U,f)\cap\omega(V,f)$. Thus
$\omega(M\cap k(\psi(M)^{<}),f)\subset\omega(M,f)\cap\omega(k(\psi(M)^{<}),f),$
and
@@
Since $\omega(M\cap k(\psi(M)^{<}),f)=\emptyset$, $M\cap k(\psi(M)^{<})=\emptyset$.
@@
[…]
Before turning to the challenges we note that the Morse nodes $q_{0}$ and $q_{7}$ identify the location of the Morse sets $M_{0}$ and $M_{2}$ of $f$, respectively, in that $M_{0}\subset|\pi^{-1}(q_{0})|$ and $M_{2}\subset|\pi^{-1}(q_{7})|$. Furthermore, as is shown in the proof of Theorem 1.2, $\mathop{\mathrm{Con}_{*}}\nolimits(M_{0})\sim\mathop{\mathrm{Con}_{*}}\nolimits(q_{0})$ and $\mathop{\mathrm{Con}_{*}}\nolimits(M_{2})\sim\mathop{\mathrm{Con}_{*}}\nolimits(q_{7})$. This represents the ideal information that can be provided by $\mathcal{G}$}.
@@
Overly refined recurrence occurs if there exists a Morse representation of $f$ that is finer than $\mathsf{MR}$}, and $\mathcal{G}$} identifies this finer Morse representation. Therefore, to satisfy the conclusion of Theorem 1.2 we collapse this information to construct the desired Morse representation.
@@
[…]
@@
By Proposition 5.3, there exists $q\in{\mathsf{MG}}({\mathcal{G}})$ such that $M\cap|\pi^{-1}(q)|\neq\emptyset$. By Proposition 5.2, $M\subset|s(k(\psi(M)))\setminus s(k(\psi(M)^{<}))|$. This implies that $\pi^{-1}(q)\cap s(k(\psi(M)))\neq\emptyset$. By Lemma 5.4, $\pi^{-1}(q)\subset s(k(\psi(M)))$. Then
$\pi^{-1}(q)\cap s(k(\psi(M)^{<}))=\emptyset$, for otherwise Lemma 5.4 implies that
$\pi^{-1}(q)\subset s(k(\psi(M)^{<}))$, contradicting the fact that $M\cap|\pi^{-1}(q)|\neq\emptyset$ and $M\subset|s(k(\psi(M)))\setminus s(k(\psi(M)^{<}))|$. Therefore, $I(M)\neq\emptyset$.
If $I(M)$ consists of a unique element, then the result is trivially true. Thus, we need to show that if $q_{0},q_{1}\in I(M)$ and $q_{0}<_{{\mathsf{MG}}({\mathcal{G}})}q^{\prime}<_{{\mathsf{MG}}({\mathcal{G}})}q_{1}$, then $q^{\prime}\in I(M)$. We provide a proof by contradiction, i.e., we assume that $q^{\prime}\not\in I(M)$.
@@
[…]
@@
Recall that $\psi:{\mathsf{MR}}\to{\mathsf{J}}^{\vee}({\mathsf{A}})$ is a poset isomorphism and $k({\mathsf{J}}^{\vee}({\mathsf{A}}))={\mathsf{J}}^{\vee}({\mathsf{N}})$. By [davey:priestley, Proposition 2.45], $X={\mathbf{1}}_{\mathsf{N}}=\bigvee_{M\in{\mathsf{MR}}}k(\psi(M))=\bigcup_{M\in{\mathsf{MR}}}k(\psi(M))$. Therefore, ${\mathcal{X}}=\bigcup_{M\in{\mathsf{MR}}}s(k(\psi(M)))$.
@@
So assume $\psi(M_{0})^{<}\not\in{\mathsf{J}}({\mathsf{A}})$. As given by Proposition 5.1 and [kalies:mischaikow:vandervorst:15, Theorem 1.2], $s$ and $k$ are isomorphisms, which implies $\psi(M_{0})^{<}\neq\emptyset$. Then by [davey:priestley, Proposition 2.45], $\psi(M_{0})^{<}$ can be written as the non-empty join $\psi(M_{0})^{<}=\bigvee_{M<M_{0}}s(k(\psi(M)))$. Thus, $\pi^{-1}(q)\cap s(k(\psi(M_{1})))\neq\emptyset$ for some $M_{1}<M_{0}$ and we repeat the argument.
[…]